In January 2017, Mitzu Co. pays $2,750,000 for a tract of land with two buildings on it. It plans to demolish Building 1 and build a new store in its place. Building 2 will be a company office; it is appraised at $690,000, with a useful life of 20 years and a $75,000 salvage value. A lighted parking lot near Building 1 has improvements (Land Improvements 1) valued at $540,000 that are expected to last another 18 years with no salvage value. Without the buildings and improvements, the tract of land is valued at $1,770,000. The company also incurs the following additional costs: Cost to demolish Building 1 $ 341,400 Cost of additional land grading 191,400 Cost to construct new building (Building 3), having a useful life of 25 years and a $400,000 salvage value 2,242,000 Cost of new land improvements (Land Improvements 2) near Building 2 having a 20-year useful life and no salvage value 173,000 Required: 1. Allocate the costs incurred by Mitzu to the appropriate columns and total each column. 2. Prepare a single journal entry to record all the incurred costs assuming they are paid in cash on January 1, 2017. 3. Using the straight-line method, prepare the December 31 adjusting entries to record depreciation for the 12 months of 2017 when these assets were in use.
Answer #1

SOLUTION

1. Allocation of costs:

Allocation of purchase price Appraised Value ($) Percent of Total Appraised Value (A) Total cost of acquisition ($) (B) Apportioned cost ($) (A*B)
Land 1,770,000 59% 2,750,000 1,622,500
Building 2 690,000 23% 2,750,000 632,500
Land Improvements 1 540,000 18% 2,750,000 495,000
Total 3,000,000 100% 2,750,000
Land Building 2 Building 3 Land Improvements 1 Land Improvements 2
Purchase price 1,622,500 632,500 495,000
Demolition 341,400
Land grading 191,400
New building (construction cost) 2,242,000
New improvements 173,000
Total 2,155,300 632,500 2,242,000 495,000 173,000

2. Journal entry:

Date Account Titles and Explanations Debit ($) Credit ($)
January 1, 2017 Land 2,155,300
Land Improvements 1 495,000
Land Improvements 2 173,000
Building 2 632,500
Building 3 2,242,000
Cash 5,697,800
(To record cost of plant assets paid in cash)

3.

Date Account Titles and Explanations Debit ($) Credit ($)
Dec.31,2017 Depreciation expense 137,705
Accumulated depreciation-Land improvement 1 27,500
Accumulated depreciation-Land improvement 2 8,650
Accumulated depreciation-Building 2 27,875
Accumulated depreciation-Building 3 73,680
(To record depreciation expense for 2017)

Accumulated depreciation-Land improvement 1 = 495,000/ 18 years = 27,500

Accumulated depreciation-Land improvement 2 = 173,000 / 20 years = 8,650

Accumulated depreciation-Building 2 = (632,500-75,000)/ 20 years = 27,875

Accumulated depreciation-Building 3 = (2,242,000 - 400,000) / 25 years = 73,680